# Membrane phononic integrated circuits

### v1

@@ X.4 Transfer Matrices
As depicted in Fig. 35(b), we can describe the change in width as a change in acoustic impedance, $Z$. This is a one-dimensional treatment analogous to how transmission lines can be modelled [55]. Acoustic impedance $Z$ for sound waves propagating in one dimension, such as in membrane waveguides, can be written as [147, 38] $Z=\rho v_{\mathrm{ph}}$, where $\rho$ is the volumetric density of the membrane and $v_{\mathrm{ph}}$ is the phase velocity of the waves. Substituting for $v_{\mathrm{ph}}$ using the standard dispersion relationship (Eq. (11)), we can write the impedance as:
@@
Here $x$ is the distance travelled, and $k$ of the acoustic wave, which depends on the waveguide width and frequency. In the case of evanescent propagation $k$ is imaginary and the amplitude exponentially decays.
Figure 35(d) illustrates that we can model a complete system involving multiple junctions and waveguides by using transfer and propagation matrices. Consider a chain of $N$ junctions in a row, describing propagation from a zeroth medium to an $N^{\mathrm{th}}$ medium. We can describe the wave amplitudes in the $i^{\mathrm{th}}$ medium as:
@@
Figure 36: Wave propagation through a two-port resonator simulated using the transfer matrix method (Eq. (125)). (a) Diagram of the simulated resonator, coupled to input and output waveguides by evanescent tunnels. Dimensions: waveguide width = $50\,\mathrm{\upmu m}$, resonator is a square with side length = $50\,\mathrm{\upmu m}$, tunnel width = $25\,\mathrm{\upmu m}$, tunnel length = $20\,\mathrm{\upmu m}$. The membrane has volumetric density $\rho=3200\,\mathrm{kg\cdot m^{-3}}$ and tensile stress $\sigma=1\,\mathrm{GPa}$, corresponding to highly stressed silicon nitride. (b) Transmission spectrum, showing a resonance at approximately $7.4\,\mathrm{MHz}$. Inset: close-up of the resonance. Yellow arrows indicate the full width half maximum of the peak, approximately $14\,\mathrm{kHz}$, corresponding to a quality factor of $Q\simeq 530$. Left and right grey boxes respectively highlight the cutoff frequency $\Omega_{c,1}$ of the tunnel and the cutoff frequency of the waveguides/resonator.
@@ X.5 Comparison of numerical methods
Figure 37: Comparison of model predictions of the coupling rate of a realistic two-port membrane resonator. (a) Illustration of the simulated device: a 50 µm square silicon nitride resonator (green) coupled to two 50 µm wide waveguides (blue) via 20 µm wide evanescent tunnels (orange) of variable length. The next two subplots show the calculated total decay rate $\gamma_{\mathrm{tot}}$ (b) and quality factor (c) of the resonator. Orange: results from calculating the overlap integral (102) using numerically calculated modeshapes (Section X.1). Blue: results from using finite element modelling to find a complex eigenfrequency (Section X.2). Green: results from using a custom finite difference time domain simulation (Section X.3). Yellow: results from using a transfer matrix model with the impedance correction in Eq. (126) (Section X.4). Note the yellow points are almost directly on top of the blue points. The coupling is calculated at the fundamental eigenfrequency of the resonator at approximately $7.62\,\mathrm{MHz}.$ Material parameters are stress $\sigma=1\,\mathrm{GPa}$ and density $\rho=3200\,\mathrm{kg\cdot m^{-3}}$.
@@
Figure 37 compares these different methods for the situation of a 2-port square membrane connected by evanescent tunnels to waveguides. This setup is the phononic version of the widely used optical Fabry-Pérot cavity [20]. In Section XI we also will show that variations on this layout correspond to a range of useful devices.
@@
There is excellent qualitative agreement between the four calculation methods. As expected we see exponential attenuation of the wave amplitude within the tunnel coupler, where the drive frequency is below the cutoff frequency and so the wavenumber is complex [22]. The exponential sensitivity of the attenuation on the tunnel length means that a wide range of quality factors can be engineered with relatively small changes in geometry.
@@
Impressively, the overlap integral method that we derived from coupled-modes theory and analytic expressions converges to the numerical methods for large tunnel lengths. The method underestimates the coupling at short tunnel lengths, where the small-coupling, perturbative approximation breaks down. At these short tunnel lengths the resonator and waveguide are sufficiently strongly coupled that their eigenmodes begin to differ significantly from the case when they are infinitely separated. The perturbative approach of CMT is more appropriate at long tunnel lengths where there is weaker coupling.
@@
The FDTD simulation achieves excellent agreement with the other numerical methods, particularly for shorter tunnels where the coupling is stronger. Performance in this regime demonstrates the flexibility of the FDTD method: because it directly solves the wave equation, it does not rely on any particular assumptions (e.g. small coupling) to hold. Indeed, the FDTD method is the best at simulating nonlinearity, strong coupling and transient behaviour over small timescales [20] (see Section X.3 above). However, in the regime of large tunnel lengths and weak coupling we also see the key flaw of FDTD methods, which is an upper limit on the observable decay rate. This limit is caused by the finite simulation finish time, which was kept constant for each of the simulations in Fig. 37.
@@
The fact that the transfer matrix method data plot a straight line of the expected slope is not surprising, as the propagation matrices in the tunnel regime are literally exponential decay operators and remain perfectly accurate with increasing distance (to the point of floating point error). What is more surprising is the nearly-perfect agreement with the finite element model, which indicates the predicted impedance mismatch is a very good estimate of the true value. This performance can be attributed to the near-field correction from microwave theory, Eq. (126). Without this correction the transfer matrix method remains qualitatively accurate but overestimates the coupling rate by $75\%$.
@@ XI Examples of engineered coupling
In addition to coupling rates there are other, intuitive insights that we can take from the CMT, such as orthogonality and symmetry. Equipped with this intuition and numerical techniques we can rapidly prototype sophisticated signal processing phononic devices. In this final section we show how, using the examples of power splitters and mode multiplexers incorporated into a single, compact resonator footprint.
@@ XI.1 Resonant arbitrary ratio power splitter
We can achieve an arbitrary splitting ratio by choosing the coupling rates. Denote the input waveguide coupling rate as $\gamma_{1}$ and the two output waveguide coupling rates as $\gamma_{2}$ and $\gamma_{3}$. A reflectionless device can be achieved by choosing $\gamma_{1}=\gamma_{2}+\gamma_{3}$, corresponding to impedance matching (see Section IX.1). In that case, the total loss or bandwidth of the resonator is $2\gamma_{1}$ (ignoring internal losses), so tuning $\gamma_{1}$ controls the bandwidth. Finally, the power splitting ratio can be set by the ratio of $\gamma_{2}$ to $\gamma_{3}$: denote the input power as $P_{1}$, the power going into waveguide 2 as $P_{2}$, and the power going into waveguide 3 as $P_{3}$. At steady state the ratio of $P_{2}$ and $P_{3}$ is related to the coupling rates as: $P_{2}/P_{3}=\gamma_{2}/\gamma_{3}$. Therefore to achieve a power splitting ratio of $P_{2}=xP_{1}$ (and $P_{3}=(1-x)P_{2}$), where $x\in[0,1]$, the coupling rates must be set to $\gamma_{2}/\gamma_{3}=P_{2}/P_{3}=x/(1-x)$.
For example, imagine we want a 50:50 impedance matched power splitter with a bandwidth of $10\,\mathrm{kHz}$. Let the coupling rates $\gamma_{\mathrm{in}}$, $\gamma_{1}$ and $\gamma_{2}$ represent the input waveguide and two output waveguide couplings. To achieve the desired bandwidth we should have $\gamma_{1}+\gamma_{2}+\gamma_{3}=2\pi\times 10\,\mathrm{kHz}$. For impedance matching we also want $\gamma_{\mathrm{in}}=\gamma_{1}+\gamma_{2}$, and for 50:50 power splitting we want $\gamma_{1}=\gamma_{2}$. Solving these equations gives that $\gamma_{\mathrm{in}}=2\gamma_{1}=2\gamma_{2}=2\pi\times 5\,\mathrm{kHz}$. Reading off of Fig. 37 using the transfer matrix or FEM method data, this suggests that the input tunnel should be approximately $28.5\,\mathrm{\mu m}$ long, and the output tunnels should be $36.3\,\mathrm{\mu m}$ long. If we wanted a different power splitting ratio of, say, 90:10, then we would just set $\gamma_{1}=9\gamma_{2}$. This corresponds to respective tunnel lengths of $L_{1}=29.7\,\mathrm{\upmu m}$ and $L_{2}=54.5\,\mathrm{\upmu m}$. Figure 37 only needs to be calculated once to be used as a reference for these different designs.
@@
Time-domain numerical simulations validate our design for the power splitter. For example, Fig. 39 shows an FDTD simulation of a three-port impedance-matched power splitter. As the simulation is initialized with waveguides and resonator at rest, the resonator must first ring-up before reaching steady-state. During this transitory period, which takes on the order of $Q$ oscillations, a gradually dimishing fraction of the incident energy is backreflected until, in the steady state, Eq. (97) is achieved and acoustic energy incident in the input flows with no back-reflections into the output ports. This is seen in Fig. 39, where the input amplitude (green trace) is initially beating as the resonator rings up, but then relaxes into a constant oscillation consistent with zero back reflection.
As seen in Fig. 37, the coupling rate is exponentially sensitive to the tunnel length, so significantly different power splitter behaviours can be engineered without compromising the device footprint. Tunnels with precisely different lengths can be realised by using a small mesh size and the top-down release fabrication method [22, 33].
@@ XI.2 Benefits and trade-offs of resonant devices
Non-resonant devices include waveguide couplers like we explored in Section VII.1. These have a high bandwidth and no transient behaviour, at the cost of a large of physical footprint. The large footprint also makes it hard to include several ports, which one can see by comparison with on-chip photonics [150, 151, 152].
Resonant devices feature multiple inputs and outputs centrally coupled to a single resonator. This design is far more compact, at the cost of requiring a ‘ring-up’ period before reaching nominal performance in the steady state. This ring up period will scale with $\gamma^{-1}=Q/\Omega$, where $\gamma$ is the total out-coupling rate of the central resonator, $Q$ is its quality factor, and $\Omega$ is the operating frequency. Larger coupling rates to input and output waveguides will produce a greater device bandwidth and shorter transient period. Conversely, bandwidth selectivity and amplitude enhancement within the resonator (if that is desirable) can be achieved by using lower coupling rates.
@@ XI.3 Resonant mode converting power splitter
Figure 40: Phononic mode converting power splitter (a) Illustration of the mode converting powersplitter. A single mode waveguide provides input to a resonant cavity. Half the energy exits on another single mode waveguide and half as the second mode in a multimode waveguie. The waveguides are coupled to the resonator by reflectionless evanescent tunnels [22] (b) Still frame from FDTD simulation. Parameters: Single mode waveguide has width 50 µm, multimode waveguide has width 100 µm. Tunnels have width 30 µm; the left tunnel has length $6\,\mathrm{\upmu m}$ and the right tunnels have length $14\,\mathrm{\upmu m}$. The resonator has dimensions $50\times 200$ µm2.
@@ XI.4 Phononic mode division multiplexer
Figure 41: Phononic mode division multiplexer. (a) Concept: the fundamental mode of the multimode waveguide drives the acoustic cavity with mirror symmetry along the $y$ axis, and therefore exits into the single mode waveguide on the right. (b) Conversely the $n=2$ waveguide mode excites the frequency-degenerate cavity mode with mirror symmetry along the $x$ axis, and therefore exits to the top waveguide. (c) FDTD simulation of the $n=1$ mode coupling to the $x$ resonant mode and out-coupling to the right waveguide. (d) FDTD simulation of the $n=2$ mode coupling to the $y$ resonant mode and out-coupling to the top waveguide. (e) FDTD simulation of a linear combination of $n=1$ and $n=2$ eigenmodes being demultiplexed. Parameters: Single mode waveguide has width 50 µm, multimode waveguide has width 100 µm. The resonator is a square with side length $100\,\mathrm{\upmu m}$. Tunnels have width 30 µm; the short tunnels have length $2.5\,\mathrm{\upmu m}$ and the long tunnels have length $5\,\mathrm{\upmu m}$.
As a final example we consider the design of a phononic mode division multiplexer. Mode division multiplexing is the technique of increasing bandwidth by encoding information in the different spatial modes of a multimode communication channel. This requires the ability to combine multiple (preferentially single-mode) inputs, encode these into different spatial modes of a multimode waveguide, and finally separate these again into single-mode outputs.
Figure 41(a-b) explains the concept for a two-mode demultiplexer. (Operated in reverse, the same device functions as a mode multiplexer.) As with the previous example, symmetry is a key technique exploited here to perform the spatial filtering. The square-shaped resonator possesses two quasi-frequency degenerate acoustic resonances, one each for $x$ and $y$ mirror symmetry. Denoting them as the $x$ and $y$ modes, we can write down coupling rates $\gamma$ between each mode and the input and output modes. There are two modes in the input waveguide, $n=1$ and $n=2$, producing input coupling rates $\gamma_{1x},\,\gamma_{1y},\,\gamma_{2x},$ and $\gamma_{2y}$, and output coupling rates $\gamma_{x\uparrow},\gamma_{y\uparrow},\gamma_{x\rightarrow}$ and $\gamma_{y\rightarrow}$. Because the fundamental mode in the input waveguide has symmetry in the $x$ axis and antisymmetry in the $y$ axis, $\gamma_{1x}\gg\gamma_{1y}$ and the $n=1$ input mode overwhelmingly excites the $x$ resonator mode. For simlar reasons the $n=2$ input mode overwhelmingly excites the $y$ resonator mode. Another set of symmetry arguments imply that the $x$ resonator mode couples to the right waveguide ($\gamma_{x\rightarrow}\gg\gamma_{x\uparrow}$) and the $y$ resonator mode couples to the top waveguide ($\gamma_{y\uparrow}\gg\gamma_{u\rightarrow}$).
@@
A variant of this design with four impedance-matched tunnel barriers coupled to single mode waveguides would function as a cross-talk free acoustic waveguide junction, analogous to the kind demonstrated in the photonic realm [154]. To maintain a single frequency of operation, care would need be taken in the fabrication process to ensure the $x$ and $y$ modes are frequency degenerate. Specifically, any difference in the mode frequencies should be less than the linewidths of the modes—for a resonator with low intrinsic dissipation, these linewidths will be primarily determined by the waveguide coupling rates. Shorter and wider tunnel couplers will produce broader linewidths, meaning greater fabrication tolerances and higher operating bandwidth. The downside will be greater crosstalk due to direct coupling between the waveguides [154].

### v2

@@ X.4 Transfer Matrices
As depicted in Fig. 35(b), we can describe the change in width as a change in acoustic impedance, $Z$. This is a one-dimensional treatment analogous to how transmission lines can be modelled [62]. Acoustic impedance $Z$ for sound waves propagating in one dimension, such as in membrane waveguides, can be written as [157, 38] $Z=\rho v_{\mathrm{ph}}$, where $\rho$ is the volumetric density of the membrane and $v_{\mathrm{ph}}$ is the phase velocity of the waves. Substituting $v_{\mathrm{ph}}$ using the standard dispersion relationship (Eq. (11)), we can write the impedance as:
@@
Here $x$ is the distance travelled and $k$ is the wavenumber, which depends on the waveguide width and frequency. In the case of evanescent propagation, $k$ is imaginary and the amplitude exponentially decays.
Figure 35(d) illustrates how we can model a complete system involving multiple junctions and waveguides by using transfer and propagation matrices. Consider a chain of $N$ junctions in a row, describing propagation from a zeroth medium to an $N^{\mathrm{th}}$ medium. We can describe the wave amplitudes in the $i^{\mathrm{th}}$ medium as:
@@
Figure 36: Wave propagation through a two-port resonator simulated using the transfer matrix method (Eq. (125)). (a) Diagram of the simulated resonator, coupled to input and output waveguides by evanescent tunnels. Dimensions: waveguide width = $50\,\mathrm{\upmu m}$, resonator is a square with side length = $50\,\mathrm{\upmu m}$, tunnel width = $25\,\mathrm{\upmu m}$, tunnel length = $20\,\mathrm{\upmu m}$. The membrane has volumetric density $\rho=3200\,\mathrm{kg\cdot m^{-3}}$ and tensile stress $\sigma=1\,\mathrm{GPa}$, corresponding to highly stressed silicon nitride. (b) Transmission spectrum, showing a resonance at approximately $7.4\,\mathrm{MHz}$. Inset: close-up of the resonance. Yellow arrows indicate the full width half maximum of the peak, approximately $14\,\mathrm{kHz}$, corresponding to a quality factor of $Q\simeq 530$, determined by the coupling losses from the resonator to both waveguides. The left grey box and right dashed line respectively highlight the cutoff frequency $\Omega_{c,1}$ of the waveguides/resonator and the tunnel.
@@ X.5 Comparison of numerical methods
Figure 37: Comparison of model predictions of the coupling rate of a realistic two-port membrane resonator. (a) Illustration of the simulated device: a 50 µm square silicon nitride resonator (green) coupled to two 50 µm wide waveguides (blue) via 20 µm wide evanescent tunnels (orange) of variable length. The next two subplots show the calculated total decay rate $\gamma_{\mathrm{tot}}$ (b) and quality factor (c) of the resonator. Orange: results from calculating the overlap integral (Eq. (102)) using numerically calculated modeshapes (Section X.1). Blue: results from using finite element modelling to find a complex eigenfrequency (Section X.2). Green: results obtained using a custom finite difference time domain simulation (Section X.3). Yellow: results from a transfer matrix model with the impedance correction (Eq. (126)) to account for varying lateral dimensions along the propagation direction. Note the yellow points are almost directly on top of the blue points. The coupling is calculated at the fundamental eigenfrequency of the resonator at approximately $7.62\,\mathrm{MHz}.$ Material parameters are stress $\sigma=1\,\mathrm{GPa}$ and density $\rho=3200\,\mathrm{kg\cdot m^{-3}}$.
@@
Figure 37 compares these different methods for the situation of a 2-port square membrane connected by evanescent tunnels to straight waveguides on either side. This setup is the phononic version of the widely used optical Fabry-Pérot cavity [20]. In Section XI we also will show that variations on this layout correspond to a range of useful devices.
@@
There is excellent qualitative agreement between the four calculation methods. As expected we see exponential attenuation of the wave amplitude within the tunnel coupler, where the drive frequency is below the cutoff frequency and the wavenumber is therefore complex [22]. The exponential sensitivity of the attenuation on the tunnel length means that a wide range of quality factors can be engineered with relatively small changes in geometry.
@@
Despite its simplifications, the overlap integral method that we derived from coupled-mode theory and analytic expressions converges to the numerical methods for large tunnel lengths. The method underestimates the coupling at short tunnel lengths, where the small-coupling, perturbative approximation breaks down. At these short tunnel lengths the resonator and waveguide are sufficiently strongly coupled that their eigenmodes begin to differ significantly from the case when they are infinitely separated. The perturbative approach of CMT is therefore naturally more appropriate at long tunnel lengths where there is weaker coupling.
@@
The FDTD simulation achieves excellent agreement with the other numerical methods, particularly for shorter tunnels where the coupling is stronger. Performance in this regime demonstrates the flexibility of the FDTD method: because it directly solves the wave equation, it does not rely on any particular assumptions (e.g. small coupling) to hold. Indeed, the FDTD method is the best at simulating nonlinearity, strong coupling and transient behaviour over small timescales [20] (see Section X.3 above). However, in the regime of large tunnel lengths and weak coupling we also see the key flaw of FDTD methods, which is a lower limit on the observable decay rate. This limit is caused by the finite simulation finish time, which was kept constant for each of the simulations in Fig. 37.
@@
The fact that the transfer matrix method data plot a straight line of the expected slope is not surprising, as the propagation matrices in the tunnel regime are just exponential decay operators and remain perfectly accurate with increasing distance (to the point of floating point error). What is more surprising is the nearly-perfect agreement with the finite element model, which indicates the predicted impedance mismatch is a very good estimate of the true value. This agreement with the other numerical techniques can be attributed to the near-field correction from microwave theory, Eq. (126). Without this correction the transfer matrix method remains qualitatively accurate but presents a systematic offset, overestimating the coupling rate by $75\%$.
@@ XI Examples of engineered coupling
In addition to coupling rates there are other, intuitive insights that we can take from the CMT, such as orthogonality and symmetry. Equipped with this intuition and numerical techniques we can rapidly prototype more sophisticated signal processing phononic devices. In this final section we show examples of this, using the example of power splitters and mode multiplexers incorporated into a single, compact resonator footprint.
@@ XI.1 Resonant arbitrary ratio power splitter
We can achieve an arbitrary splitting ratio by precisely choosing the coupling rates. Denote the input waveguide coupling rate as $\gamma_{1}$ and the two output waveguide coupling rates as $\gamma_{2}$ and $\gamma_{3}$. A reflectionless device can be achieved by choosing $\gamma_{1}=\gamma_{2}+\gamma_{3}$, corresponding to impedance matching (see Section IX.1). In that case, the total loss or bandwidth of the resonator is $2\gamma_{1}$ (ignoring internal losses), so tuning $\gamma_{1}$ controls the bandwidth. Finally, the power splitting ratio can be set by the ratio of $\gamma_{2}$ to $\gamma_{3}$. Indeed, let us denote denote the input power as $P_{1}$, the power going into waveguide 2 as $P_{2}$, and the power going into waveguide 3 as $P_{3}$. At steady state the ratio of $P_{2}$ and $P_{3}$ is related to the coupling rates as: $P_{2}/P_{3}=\gamma_{2}/\gamma_{3}$. Therefore to achieve a power splitting ratio of $P_{2}=xP_{1}$ (and $P_{3}=(1-x)P_{2}$), where $x\in[0,1]$, the coupling rates must be set to $\gamma_{2}/\gamma_{3}=P_{2}/P_{3}=x/(1-x)$.
For example, consider desiring a 50:50 impedance-matched power splitter with a frequency bandwidth of $10\,\mathrm{kHz}$. Let the coupling rates $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$ represent the input waveguide and two output waveguide couplings. To achieve the desired bandwidth we should have $\gamma_{1}+\gamma_{2}+\gamma_{3}=2\pi\times 10\,\mathrm{kHz}$. For impedance matching we also want $\gamma_{1}=\gamma_{2}+\gamma_{3}$, and for 50:50 power splitting we want $\gamma_{2}=\gamma_{3}$. Solving these equations gives that $\gamma_{1}=2\gamma_{2}=2\gamma_{3}=2\pi\times 5\,\mathrm{kHz}$. Reading off of Fig. 37 using the transfer matrix or FEM method data, this suggests that the input tunnel should be approximately $15\,\mathrm{\mu m}$ long, and the output tunnels should be $18\,\mathrm{\mu m}$ long. If we desire a different power splitting ratio of 90:10 we would set $\gamma_{2}=9\gamma_{3}$. This corresponds to respective tunnel lengths of $L_{2}=15\,\mathrm{\upmu m}$ and $L_{3}=25\,\mathrm{\upmu m}$. Figure 37 only needs to be calculated once to be used as a reference for these different designs.
@@
Time-domain numerical simulations validate our design for the power splitter. For example, Fig. 39 shows an FDTD simulation of a three-port impedance-matched power splitter. As the simulation is initialized with waveguides and resonator at rest, the resonator must first ring-up before reaching steady-state. During this transitory period, which takes on the order of $Q$ oscillations, a gradually dimishing fraction of the incident energy is backreflected until, in the steady state, acoustic energy incident in the input port flows with no back-reflections into the output ports. This is seen in Fig. 39, where the input amplitude (green trace) is initially beating as the resonator rings up, but then relaxes into a constant oscillation consistent with zero back reflection.
As seen in Fig. 37, the coupling rate is exponentially sensitive to the tunnel length, so significantly different power splitter behaviours can be engineered without compromising the device footprint. Coupling rates also strongly depend on the tunnel width, providing an additional independent means of control. Tunnels with precisely different lengths or widths can be realised by using a small mesh size and the top-down release fabrication method [22, 33].
@@ XI.2 Benefits and trade-offs of resonant devices
Non-resonant devices include waveguide couplers like we explored in Section VII.1. These have a high bandwidth and no transient behaviour, at the cost of a large physical footprint. The large footprint also makes it hard to include several ports, which one can see by comparison with on-chip photonics [160, 161, 162].
Resonant devices feature multiple inputs and outputs centrally coupled to a single resonator. This design is far more compact, at the cost of requiring a ‘ring-up’ period before reaching nominal performance in the steady state. This ring up period will scale with $\gamma^{-1}=Q/\Omega$, where $\gamma$ is the total out-coupling rate of the central resonator, $Q$ is its quality factor, and $\Omega$ is the operating frequency. Larger coupling rates to input and output waveguides will produce a greater device bandwidth and shorter transient period. Conversely, bandwidth selectivity and amplitude enhancement within the resonator (if desirable) can be achieved by using lower coupling rates.
@@ XI.3 Resonant mode converting power splitter
Figure 40: Phononic mode converting power splitter (a) Illustration of the mode converting powersplitter. A single mode waveguide provides input to a resonant cavity. One third of the incoming power exits on another single mode waveguide and the other two thirds as the second mode in a multimode waveguide. The waveguides are coupled to the resonator by reflectionless evanescent tunnels [22] (b) Still frame from FDTD simulation. Parameters: Single mode waveguide has width 50 µm, multimode waveguide has width 100 µm. Tunnels have width 30 µm; the left tunnel has length $6\,\mathrm{\upmu m}$ and the right tunnels have length $14\,\mathrm{\upmu m}$. The resonator has dimensions $50\times 200$ µm2.
@@ XI.4 Phononic mode division multiplexer
Figure 41: Phononic mode division multiplexer. (a) Concept: the $n=1$ mode of the multimode waveguide drives the acoustic cavity mode with mirror symmetry along the $x$ axis, and therefore exits into the right single mode output waveguide because $\gamma_{x\rightarrow}\gg\gamma_{x\uparrow}$. (b) Conversely the $n=2$ waveguide mode excites the frequency-degenerate cavity mode with mirror symmetry along the $y$ axis, and therefore exits to the top single mode output waveguide because $\gamma_{y\uparrow}\gg\gamma_{y\rightarrow}$. (c) FDTD simulation of the $n=1$ mode coupling to the $x$ resonant mode and out-coupling to the right waveguide. (d) FDTD simulation of the $n=2$ mode coupling to the $y$ resonant mode and out-coupling to the top waveguide. (e) FDTD simulation of a linear combination of $n=1$ and $n=2$ eigenmodes being demultiplexed. Parameters: Single mode waveguide has width 50 µm, multimode waveguide has width 100 µm. The resonator is a square with side length $100\,\mathrm{\upmu m}$. Tunnels have width 30 µm; the short tunnels have length $2.5\,\mathrm{\upmu m}$ and the long tunnels have length $5\,\mathrm{\upmu m}$. (f) FDTD simulation of a four-way acoustic waveguide junction with minimal crosstalk. Parameters: central square resonator has side length 50 µm, the tunnels have width 20 µm and length 5 µm, and the waveguides have width 30 µm.
As a final example we consider the design of an acoustic mode division multiplexer. Mode division multiplexing is a technique to increase the bandwidth by encoding information in the different spatial modes of a multimode communication channel [164]. This requires the ability to combine multiple (preferentially single-mode) inputs, encode these into different spatial modes of a multimode waveguide, and finally separate these again into single-mode outputs.
Figure 41(a-b) explains the concept for a two-mode demultiplexer. (Operated in reverse, the same device functions as a mode multiplexer.) As with the previous example, symmetry is a key technique exploited here to perform the spatial filtering. The square-shaped resonator possesses two quasi-frequency degenerate acoustic resonances, one each for $x$ and $y$ mirror symmetry. Denoting them as the $x$ and $y$ modes, we can write down coupling rates $\gamma$ between each mode and the input and output modes. There are two modes in the input waveguide, $n=1$ and $n=2$, producing input coupling rates $\gamma_{1x},\,\gamma_{1y},\,\gamma_{2x},$ and $\gamma_{2y}$, and output coupling rates $\gamma_{x\uparrow},\gamma_{y\uparrow},\gamma_{x\rightarrow}$ and $\gamma_{y\rightarrow}$. Because the $n=1$ mode in the input waveguide has symmetry in the $x$ axis and antisymmetry in the $y$ axis, $\gamma_{1x}\gg\gamma_{1y}$ and it overwhelmingly couples to the $x$ resonator mode. For similar reasons the $n=2$ input mode overwhelmingly couples to the $y$ resonator mode. Another set of symmetry arguments imply that the $x$ resonator mode couples to the right output single mode waveguide ($\gamma_{x\rightarrow}\gg\gamma_{x\uparrow}$) and the $y$ resonator mode couples to the top output single mode waveguide ($\gamma_{y\uparrow}\gg\gamma_{u\rightarrow}$).
@@
A variant of this design with four impedance-matched tunnel barriers coupled to single mode waveguides would function as a cross-talk free acoustic waveguide junction (Fig. 41(f)), analogous to the kind demonstrated in the photonic realm [165]. To maintain a single frequency of operation, care would need be taken in the fabrication process to ensure the $x$ and $y$ modes are frequency degenerate. Specifically, any difference in the mode frequencies should be less than the linewidths of the modes—for a resonator with low intrinsic dissipation, these linewidths will be primarily determined by the waveguide coupling rates. Shorter and wider tunnel couplers will produce broader linewidths, meaning greater fabrication tolerances and higher operating bandwidth. The downside will be greater crosstalk due to direct coupling between the waveguides (against the symmetry restrictions), due to overlap of the evanescent fields of the waveguide modes [165].